Find the nth term formula for the geometric sequence whose first term is 18 and whose second term is 9.

Answer by stanbon(48558)

Find the nth term formula for the geometric sequence whose first term is 18 and whose second term is 9.
nth term formula a(n)=ar^(n-1)
a=18
r=9/18=1/2
nth term is a(n)=18(1/2)^(n-1)
Cheers,
Stan H.